One number is five greater than another. The product of the numbers is 36. find the numbers.
Answer by shree840(260)

let the other number be x
one number is 5 more than other =x+5
product=x(x+5)=36
x^2+5x=36
x^2+5x-36=0
factoring
x^2+9x-4x-36=0
x(x+9)-4(x+9)=0
(x-4)(x+9)=0
x-4=0 or x=4 other number is 5 more =5+4=9
Two numbers are 4,9